I'm having trouble solving the equation: (x -7):(x+7) = 2:9.

Found 2 solutions by prabhjyot, JimMarshall:
Answer by prabhjyot(165)

(x -7):(x+7) = 2:9.
This equation can be written as
(x-7)/(x+7)=2/9
Now cross multiply , we get
2(x+7)= 9(x-7)
2x+ 14 =9x -63
Combining like terms
2x- 9x= -63 -14
- 7x= -77
x= -77/-7
x= 11

Answer by JimMarshall(82)

==>x -7):(x+7) = 2:9.
This equation can be written as
==>(x-7)/(x+7)=2/9
==>2(x+7)= 9(x-7)
==>2x+ 14 =9x -63
==>2x- 9x= -63 -14
==>- 7x= -77
==>x= -77/-7
==>x= 11
Ans